Alice travels 4 times faster than Gloria. Traveling in opposite directions, they are 280 miles apart after 8 hours. Find their rates of travel.
Answer by TimothyLamb(4379)

x = alice speed
y = gloria speed
x = 4y
---
s = d/t
---
their relative speed:
x + y = 280/8
x + y = 35
4y + y = 35
5y = 35
y = 7
x = 4y
x = 4*7
x = 28
---
answer:
x = alice speed = 28 mph
y = gloria speed = 7 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php